(10 points) Solve the following differential equation: ydr + edy 1-zºy2 + udt = 0. = constant.
Answer #1

+ x dx = 0 y dx tx dy 1-x²y2 d (y) I - (vyje dcxy) 1 - fly)? +7 dx = 0 + x dx x dx) = 0 day) T +/x dx = 0 dx 1-X² itxy 1-ry x